OpenStudy (anonymous):

find an equation for tangent line to y=f(x) at the specified point. f(x)=xe^(-x) where x=0

OpenStudy (anonymous):

tangent is dy/dx

OpenStudy (anonymous):

what are the step into finding the answer?

OpenStudy (anonymous):

take the derivative of xe^(-x) and then put x=0

OpenStudy (anonymous):

derivative =-e^(-x) (-1+x) =-e^(-0)(-1+(0)) =1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the equation for tangent line is y=x

OpenStudy (anonymous):

why its y=x? you found 1

OpenStudy (anonymous):

the book answer is y=x, im guessing where y=mx+b so y=1(x)

OpenStudy (anonymous):

yes its y-a=mx+b and you are taking the tangent at (0,0) so a,b=0 and the slope m=1 so its y=x sorry for answer before

OpenStudy (anonymous):

thanks for the help